OpenStudy (anonymous):

Given f(x)= x-5/3, solve for f^-1 (3)

OpenStudy (anonymous):

To find the inverse: Replace f(x) with y Switch x's and y's, so put x where y is and x where y is. Solve for y Replace y with f^-1(x)

OpenStudy (anonymous):

Also, which of the following is the conjugate of a complex number with 2 as the real part and -8 as the imaginary part?

OpenStudy (anonymous):

very confused, let me read that

OpenStudy (anonymous):

I did that and couldn't get the answer @iambatman

OpenStudy (anonymous):

Show me.

OpenStudy (anonymous):

x=y-5/3

OpenStudy (anonymous):

i honestly don't know

OpenStudy (anonymous):

Let's follow the steps, and we'll worry about f^-1(3) after.

OpenStudy (the_fizicx99):

\(\ \sf \Large f(x) = \dfrac{x - 5}{3} \) 1. replace f(x) with y 2. Switch x where y is and x where y is. 3. Solve for y again. x = \(\ \sf \dfrac{y - 5}{3} \) 3x = y - 5 3x + 5 = y Replace the "y" you just solved for with \(\ f^{-1}(x) \) And that's your inverse function, \(\ \sf f^{-1}(x) = 3x + 5 \) Now plug in 3 for "x" and that'll be f^-1(3)

OpenStudy (anonymous):

f(x) = x-5/3 y = x-5/3 x = y - 5/3 You got up to that point, so nice. Now solve for y. x+5/3 = y f^-1(x) = x+5/3 Now just plug in the 3 f^-1(3) = 3+5/3 You can do that at least :)

OpenStudy (anonymous):

Oh is it x-5/3 or (x-5)/3 >_>

OpenStudy (anonymous):

Either way you have answer from both sides, ty fizi :3

OpenStudy (anonymous):

I got 14

OpenStudy (anonymous):

Thankyou